The Modern Grocery has cashews that sell for $3.50 a pound and peanuts that sell for $2.25 a pound. How much of each must Albert, the grocer, mix to get 50 pounds of mixture that he can sell for $3.00 per pound. Express the problem as a system of linear equations and solve using the method of your choice to find the solution of the problem. (Identify your variables and show all work.)
I tried: PRICE 3.5c + 2.25p = 150
AMT c + p = 50
c- cashews p-peanuts
I got p- 9.09 lbs
C- 40.91 lbs

Answer by stanbon(75887)

The Modern Grocery has cashews that sell for $3.50 a pound and peanuts that sell for $2.25 a pound. How much of each must Albert, the grocer, mix to get 50 pounds of mixture that he can sell for $3.00 per pound. Express the problem as a system of linear equations and solve using the method of your choice to find the solution of the problem. (Identify your variables and show all work.)
I tried: PRICE 3.5c + 2.25p = 150
AMT c + p = 50
c=50-p
3.5(50-p)+2.25p = 150
175-3.5p+2.25p=150
-1.25p=-25
p=20
c=30
Cheers,
Stan H.